# A Modified 3D Biconical Outflow Model: Spatial Constraints on AGN-driven Outflows

### v1

@@ 5.4 Outflow Energetics
We estimate mass outflow rates ($\dot{\mathrm{M}}_{\mathrm{out}}$), momentum rates ($\dot{\mathrm{P}}_{\mathrm{out}}$), and kinetic powers ($\dot{\mathrm{E}}_{\mathrm{out}}$) of local type 2 AGNs to investigate feedback efficiency. These outflow rates can be determined for conical outflows as (e.g., Maiolino12; Fiore17)
$\dot{\mathrm{M}}_{\mathrm{out}}=3\;\mathrm{M}_{\mathrm{out}}\frac{v_{\mathrm{out}}}{\mathrm{R}_{\mathrm{out}}}$ (9)
$\dot{\mathrm{P}}_{\mathrm{out}}=\dot{\mathrm{M}}_{\mathrm{out}}v_{\mathrm{out}}$ (10)
$\dot{\mathrm{E}}_{\mathrm{out}}=\frac{1}{2}\dot{\mathrm{M}}_{\mathrm{out}}v^{2}_{\mathrm{out}}$ (11)
where $\mathrm{M}_{\mathrm{out}}$, $v_{\mathrm{out}}$, and $\mathrm{R}_{\mathrm{out}}$ are ionized gas mass, outflow velocity, and outflow size, respectively. We estimate ionized gas mass from the extinction-corrected [O III] luminosity, using Eq. 29 in Veilleux20 and a fixed electron density of 1000 cm-3.
@@
In the case of outflow velocity, various definitions were adopted by previous studies, i.e., W80 (e.g., Wylezalek20; Ruschel-Dutra21), $\sqrt{v^{2}_{\mathrm{[OIII]}}+\sigma^{2}_{\mathrm{[OIII]}}}$ (e.g., Rakshit18; Kim23), or $v_{\mathrm{broad}}+2\sigma_{\mathrm{broad}}$ (e.g., Rupke13; Fiore17) to represent the flux-weighted mean outflow velocity. We derive the mass-weighted mean outflow velocity, using the radial profiles of outflow velocity $v(R)=V_{max}(1-R/R_{out})$ and flux $f(R)=f_{0}e^{-5R/R_{out}}$ in our model and assuming an electron density profile, $n_{e}(R)$ $\propto$ $1/R^{2}$. Since the mass is proportional to L/[$_{\mathrm{[OIII]}}/n_{e}$, the mass-weighted mean velocity is derived as:
$\begin{split}v_{\mathrm{out}}&=\frac{\int^{R_{out}}_{0}v(R)\mathrm{M}_{\mathrm{out}}(R)\,R^{2}dR}{\int^{R_{out}}_{0}\mathrm{M}_{\mathrm{out}}(R)\,R^{2}dR}\\ &=\frac{\int^{R_{out}}_{0}V_{max}(1-\frac{R}{R_{out}})\times f_{0}e^{-\frac{5R}{R_{out}}}n^{-1}_{e}(R)R^{2}\,dR}{\int^{R_{out}}_{0}f_{0}e^{-\frac{5R}{R_{out}}}n^{-1}_{e}(R)R^{2}\,dR}\\ &=V_{max}\frac{\int^{1}_{0}t^{4}(1-t)e^{-5t}\,dt}{\int^{1}_{0}t^{4}e^{-5t}\,dt}\simeq 0.31\;V_{max}\end{split}$ (12)
@@
The mass-weighted mean outflow velocity is obtained from the maximum velocity as $v_{\mathrm{out}}=0.31V_{max}$. In addition, since the momentum rate and kinetic power are proportional to $v_{\mathrm{out}}^{2}$ and $v_{\mathrm{out}}^{3}$, respectively, we calculate the mass-weighted means of $v^{2}$ and $v^{3}$ as $\sim 0.14\;V^{2}_{max}$ and $\sim 0.07\;V^{3}_{max}$. These relations are then applied to calculate the three outflow rate parameters.
@@
Note that the launching velocity cannot be directly measured from observational data. Instead, we test potential proxies of $V_{max}$ that can be calculated from the observed V[OIII] and $\sigma_{\mathrm{[OIII]}}$, using the mock SDSS data (Section 4.2). We find that two proxies show good agreement with the input $V_{max}$. One is a quadratic sum of $|\mathrm{V}_{\mathrm{[OIII]}}|$ and 3 $\sigma_{\mathrm{[OIII]}}$, which is similar to Bae16 or RodriguezZaurin13, and the other is a direct summation ($|\mathrm{V}_{\mathrm{[OIII]}}|$ + 3 $\sigma_{\mathrm{[OIII]}}$), which is similar to Rupke13 or Fiore17. In Figure 14, we compare the two proxies, showing that both reproduce the launching velocity with a scatter of 0.2 dex. Here, we use the first proxy (i.e., quadratic sum).
@@
Lastly, we utilize two outflow sizes as upper and lower limits. As discussed in Section 5.3, we first adopt our $\sigma_{\mathrm{[OIII]}}$-based sizes calculated from the kinematic outflow size–luminosity relation as upper limits and determine the lower limits of outflow rates. Second, we utilize the NLR sizes calculated from the NLR size–luminosity relation, which was derived from nearby AGNs (Polack24). We determine lower limits of the outflow sizes by dividing the NLR sizes by three, which is a typical Rout/R[OIII] ratio from Fischer18; Polack24, and calculate upper limits of the outflow rates. Note that these outflow sizes are typically 0.4 - 0.6 dex smaller than the $\sigma_{\mathrm{[OIII]}}$-based sizes within the dynamic range of the [O III] luminosity.
@@
Figure 15: Mass outflow rate (log $\dot{\mathrm{M}}_{\mathrm{out}}$, top), momentum rate (log $\dot{\mathrm{P}}_{\mathrm{out}}$, middle), and kinetic power (log $\dot{\mathrm{E}}_{\mathrm{out}}$, bottom) as a function of bolometric luminosity of the SDSS type 2 AGNs. The red and blue contours show the energetics based on the upper limits (red contours) and lower limits (blue contours) of the outflow sizes. The energetics of individual AGNs from the literature are denoted with various symbols.
In Figure 15, we compare the upper and lower limits of the outflow rates with luminosity to investigate whether the outflows are energetic enough for efficient feedback. Despite the large scatter, the estimated outflow rates (contours) fall within the ranges of $\dot{\mathrm{M}}_{\mathrm{out}}\sim 1-100\;\mathrm{L}_{\mathrm{bol}}/\eta c^{2}\;(\eta=0.1)$, $\dot{\mathrm{P}}_{\mathrm{out}}\sim 0.01-1\;\mathrm{L}_{\mathrm{bol}}/c$, and $\dot{\mathrm{E}}_{\mathrm{out}}\sim 10^{-6}-10^{-3}\;\mathrm{L}_{\mathrm{bol}}$. These ranges are consistent with previous studies, except for some discrepancies due to different definitions of outflow parameters. For example, Fiore17 sample, which has kinetic powers mostly in the range of 0.01 - 1% of Lbol, has much higher $v_{\mathrm{out}}$ values since it was defined based on the kinematics of the broad component ($|v_{\mathrm{broad}}|+2\sigma_{\mathrm{broad}}$). In contrast, we compute $v_{\mathrm{out}}$ using the first and second moments of the total [O III] profile, leading to smaller $v_{\mathrm{out}}$ at a given luminosity. This difference becomes more evident in $\dot{\mathrm{E}}_{\mathrm{out}}$ since it is proportional to $v^{3}_{\mathrm{out}}$.
Theoretical studies have suggested thresholds of kinetic power relative to bolometric luminosity (i.e., $\dot{\mathrm{E}}_{\mathrm{out}}/\mathrm{L}_{\mathrm{bol}}\sim 0.5-5\%$, e.g., Hopkins10a; Zubovas12; King15), that can suppress star formation. The estimated kinetic power is mostly much lower than the thresholds, implying that feedback may not be efficient. However, it is important to note that these outflow rates are based only on ionized outflows, which contribute a relatively low fraction of the total outflow mass (e.g., Fluetsch19). Nevertheless, 0.85% or 0.14% of the sample has $\dot{\mathrm{E}}_{\mathrm{out}}/\mathrm{L}_{\mathrm{bol}}$ higher than 0.5% when using outflow sizes based on Polack24 or Kim23 relations, respectively. These AGNs with extremely high kinetic powers also exhibit high outflow velocities, making them excellent targets for spatially resolved outflow studies in detail.
5.5 Limitations of the current model
@@
Despite the modifications, there are still limitations in the model and the mock data analysis. First, since we adopt analytic functions for the flux and velocity distributions of outflows, our model cannot describe the detailed structures in outflows. For example, various studies of nearby AGNs found clumpy regions within the bicone structure (e.g., Venturi18; Mingozzi19; Kakkad23).
@@
Second, our model simulates no flux errors in the mock data. For example, the $\sigma_{\mathrm{[OIII]}}$-based method adopts the radius of the furthest [O III]-detected pixel as an outflow size when the $\sigma_{\mathrm{[OIII]}}/\sigma_{*}$ does not converge to unity. In this case, the S/N becomes critical for determining the outflow size. Also, we did not test spectroastrometry-based methods (e.g., Carniani15; Singha22), which define the outflow size as a spatial offset between the center and flux centroid of the [O III] emission. Reasonable errors are necessary to fit the 2D flux map when estimating the flux centroid.
@@
Third, we assumed that all PDFs in our MC simulations are independent since we do not have a well-described dependence among them. However, since some of the input parameters are likely to be correlated, the PDFs should be described as multivariate. Previous studies investigated the correlations between outflow parameters and bolometric luminosity (or Eddington ratio). The opening angle and outflow velocity are also suggested to have positive (Bae16), negative (Muller-Sanchez11), or no correlations (Fischer13). Therefore, MC simulations with a proper treatment of correlations between outflow parameters may better reproduce the observed gas kinematics.
6 Summary
@@
This work has been supported by the Basic Science Research Program through the National Research Foundation of Korean Government (2021R1A2C3008486).

### v2

@@ V.4 Outflow Energetics
In this section, we present estimates of the mass outflow rate ($\dot{\mathrm{M}}_{\mathrm{out}}$), momentum rate ($\dot{\mathrm{P}}_{\mathrm{out}}$), and kinetic power ($\dot{\mathrm{E}}_{\mathrm{out}}$) for local type 2 AGNs to assess the efficiency of AGN feedback. For conical outflows, these quantities are computed as (e.g., Maiolino et al., 2012; Fiore et al., 2017)
$\dot{\mathrm{M}}_{\mathrm{out}}=3\,\mathrm{M}_{\mathrm{out}}\frac{v_{\mathrm{out}}}{\mathrm{R}_{\mathrm{out}}},$ (9)
$\dot{\mathrm{P}}_{\mathrm{out}}=\dot{\mathrm{M}}_{\mathrm{out}}\,v_{\mathrm{out}},$ (10)
$\dot{\mathrm{E}}_{\mathrm{out}}=\frac{1}{2}\,\dot{\mathrm{M}}_{\mathrm{out}}\,v_{\mathrm{out}}^{2},$ (11)
[…]
@@
A variety of velocity definitions have been used in previous studies to approximate the characteristic outflow velocity, including W80 (e.g., Wylezalek et al., 2020; Ruschel-Dutra et al., 2021), $\sqrt{v_{\rm[OIII]}^{2}+\sigma_{\rm[OIII]}^{2}}$ (e.g., Rakshit and Woo, 2018; Kim et al., 2023), and $v_{\rm broad}+2\sigma_{\rm broad}$ (e.g., Rupke and Veilleux, 2013; Fiore et al., 2017). In our analysis, we derive a mass-weighted mean outflow velocity based on the radial profiles of the intrinsic velocity ($v(R)=V_{\rm max}(1-R/R_{\rm out})$) and flux ($f(R)=f_{0}e^{-5R/R_{\rm out}}$) in our model, assuming an electron density profile $n_{e}(R)\propto R^{-2}$. Since the mass is proportional to L/[$_{\mathrm{[OIII]}}/n_{e}$, the mass-weighted mean velocity becomes
$\begin{split}v_{\mathrm{out}}&=\frac{\int_{0}^{R_{\rm out}}v(R)\,\mathrm{M}_{\mathrm{out}}(R)\,R^{2}\,dR}{\int_{0}^{R_{\rm out}}\mathrm{M}_{\mathrm{out}}(R)\,R^{2}\,dR}\\ &=\frac{\int_{0}^{R_{\rm out}}V_{\rm max}\!\left(1-\frac{R}{R_{\rm out}}\right)f_{0}e^{-5R/R_{\rm out}}n_{e}^{-1}(R)\,R^{2}\,dR}{\int_{0}^{R_{\rm out}}f_{0}e^{-5R/R_{\rm out}}n_{e}^{-1}(R)\,R^{2}\,dR}\\ &=V_{\rm max}\,\frac{\int_{0}^{1}t^{4}(1-t)e^{-5t}\,dt}{\int_{0}^{1}t^{4}e^{-5t}\,dt}\simeq 0.31\,V_{\rm max}.\end{split}$ (12)
@@
Thus, the mass-weighted mean outflow velocity is approximately $v_{\mathrm{out}}=0.31\,V_{\rm max}$. Since the momentum and kinetic power scale with $v_{\mathrm{out}}^{2}$ and $v_{\mathrm{out}}^{3}$, respectively, their mass-weighted mean values correspond to $\sim 0.14\,V_{\rm max}^{2}$ and $\sim 0.07\,V_{\rm max}^{3}$. These relations are used to calculate $\dot{\mathrm{M}}_{\mathrm{out}}$, $\dot{\mathrm{P}}_{\mathrm{out}}$, and $\dot{\mathrm{E}}_{\mathrm{out}}$.
@@
Note that the launching velocity cannot be directly measured from observational data. Instead, we test several potential proxies for $V_{\rm max}$ that can be computed from the observed V[OIII] and $\sigma_{\rm[OIII]}$, using the mock SDSS sample (Section IV.2). We find that two commonly used prescriptions show good agreement with the input $V_{\rm max}$. The first is the quadratic sum of $|\mathrm{V}_{\rm[OIII]}|$ and $3\,\sigma_{\rm[OIII]}$, similar to that adopted by Bae and Woo (2016) and Rodríguez Zaurín et al. (2013). The second is the direct sum $|\mathrm{V}_{\rm[OIII]}|+3\,\sigma_{\rm[OIII]}$, following Rupke and Veilleux (2013) and Fiore et al. (2017). Figure 14 demonstrates that both proxies recover the launching velocity with a scatter of $\sim$0.2 dex. In the following analysis, we adopt the first proxy (i.e., the quadratic sum).
@@
Lastly, we use two different outflow size estimates to set upper and lower limits on the energetics. First, we adopt the $\sigma_{\rm[OIII]}$-based sizes inferred from the kinematic outflow size–luminosity relation as upper limits (see Section V.3), and compute the corresponding lower limits on the outflow rates. Second, we employ the NLR sizes from the NLR size–luminosity relation derived for nearby AGNs (Polack et al., 2024). We convert these NLR sizes to outflow sizes by dividing by three, following the typical ratio of $\mathrm{R_{out}}/\mathrm{R_{[OIII]}}$ reported in Fischer et al. (2018) and Polack et al. (2024), and then calculate the corresponding upper limits on the outflow rates. These NLR-based sizes are typically 0.4–0.6 dex smaller than the $\sigma_{\rm[OIII]}$-based sizes in the dynamic range of the [O III] luminosity.
@@
Figure 15: Mass outflow rate (log $\dot{\mathrm{M}}_{\mathrm{out}}$, top), momentum rate (log $\dot{\mathrm{P}}_{\mathrm{out}}$, middle), and kinetic power (log $\dot{\mathrm{E}}_{\mathrm{out}}$, bottom) as a function of the bolometric luminosity of the SDSS type 2 AGNs. The red and blue contours show the energetics based on the upper limits (red contours) and lower limits (blue contours) of the outflow sizes. The energetics of individual AGNs from the literature are denoted with various symbols.
[…]
V.5 Limitations of the current model
@@
Despite these improvements, our model and mock data analysis still have several limitations. First, because we adopt analytic functions for the flux and velocity distributions, the model cannot capture the detailed substructures of real outflows. High-resolution studies of nearby AGNs have revealed clumpy and irregular features within the bicone (e.g., Venturi et al., 2018; Mingozzi et al., 2019; Kakkad et al., 2023), which are not represented in our simplified prescription.
@@
Second, our mock data do not include flux uncertainties. For instance, the $\sigma_{\rm[OIII]}$-based method defines the outflow size as the radius of the outermost [O III]-detected pixel when the $\sigma_{\mathrm{[OIII]}}/\sigma_{*}$ does not converge to unity; in such cases, the S/N becomes critical in determining the size. We also did not test spectroastrometry-based techniques (e.g., Carniani et al., 2015; Singha et al., 2022), which estimate the outflow size from the spatial offset between the center and the flux centroid of the [O III] emission. These methods require realistic flux errors to properly fit the 2D flux distribution and determine the centroid.
@@
Third, we assume that all parameter PDFs in the MC simulations are independent, owing to the lack of well-established relations among them. In reality, several outflow parameters are expected to be correlated, implying that multivariate PDFs may be more appropriate. Previous studies have reported correlations between outflow properties and bolometric luminosity (or Eddington ratio). The opening angle and outflow velocity are also suggested to have positive (Bae and Woo, 2016), negative (Müller-Sánchez et al., 2011), or no correlations (Fischer et al., 2013). Future MC simulations incorporating such parameter correlations may better reproduce the observed gas kinematics.
VI Summary
@@
This work has been supported by the Basic Science Research Program through the National Research Foundation of the Korean Government (2021R1A2C3008486).